OpenStudy (brooke1556):

Simplify 9 over quantity of 3 minus 2 i.

OpenStudy (adonine):

Subtract and simplify: The quantity x squared over the binomial x plus 3, end quantity, minus the quantity 9 over the binomial x plus 3, end quantity.

OpenStudy (brooke1556):

\[\frac{ 9 }{ 3 - 2i }\]

jhonyy9 (jhonyy9):

multiplie the numerator and denominator by conjugate of 3-2i do you know what mean conjugate of a complex number ?

OpenStudy (lapinblanc):

\[\frac{ 9 }{ 3-2i }\times \frac{3+2i}{3+2i}\] Because 3+2i is your conjugate and you want to divide it by itself so that it simplifies to 1 and it does not change the equation. Remember \[i ^{2}=-1\] Simplify

OpenStudy (lapinblanc):

I mention the simplifying of the conjugate so that you know that you are not changing the equation. Do not simplify it before you multiply it to your original equation. Multiply to equation, then simplify.